OpenStudy (imterribleatmath):

PLEASE HELP WILL GIVE MEDAL . given that (-9,3) is on the graph of f(x) find the corresponding point for the function f(-3x)

OpenStudy (tkhunny):

Have you tried an example? f(x) = ax f(-9) = -9a = 3 ==> a = 3/-9 = -1/3 ==> f(x) = (-1/3)x f(3x) = (-1/3)(3x) = -x Is it the same for ALL such functions, given only the one point? g(x) = x + b g(-9) = -9 + b = 3 ==> b = 12 ==> g(x) = x + 12 g(3x) = 3x + 12 Looks to me like you have some variation on your hands.

OpenStudy (imterribleatmath):

i have no clue what i'm doing

satellite73 (satellite73):

you don't have very much info, so this should be relatively simple

OpenStudy (imterribleatmath):

i'm horrible at math and could use some help

OpenStudy (imterribleatmath):

@hayhayz

OpenStudy (imterribleatmath):

@johnweldon1993

OpenStudy (agent0smith):

You have the point (-9,3) which means f(-9) = 3 the corresponding point means you need to find the x-value so that: f(-3x) = 3 Any idea...?

OpenStudy (imterribleatmath):

i still don't really get it, But would f(-3x)=3 be the answer @agent0smith

OpenStudy (agent0smith):

No. f(-9) = 3 f(-3x) = 3 Stare at both of those for a few seconds. See if you can figure out what x needs to be, for them to be the exact same thing.

OpenStudy (imterribleatmath):

@agent0smith

OpenStudy (imterribleatmath):

would it be 3,3 ?

OpenStudy (tgstudios):

x = 3 that is the only way it would work

OpenStudy (imterribleatmath):

@agent0smith @TGstudios

OpenStudy (imterribleatmath):

iv sorta got it down, how do i know whether the point is on the y or x axis

OpenStudy (imterribleatmath):

how do i know whether it goes in the y or x direction

OpenStudy (agent0smith):

Your last two questions are entirely unrelated to this question. If a point is on the x axis, it'll look like (x, 0) (ie the y value will be 0) If a point is on the y axis, it'll look like (0, y) (ie the x value will be 0) "how do i know whether it goes in the y or x direction" I don't really know what you're asking.